OpenStudy (anonymous):

HELP:::Multiply and simplify 2/3 x 2 1/5 x 3/4 A. 1/5 B.3/5 C.1 1/20 D.1 1/10

OpenStudy (paki):

2/3 x 2 1/5 x 3/4 = 2/3*11/5*3/4 = 66/60 = 11/10

OpenStudy (anonymous):

@paki THANKS

OpenStudy (paki):

pleasure :-)